# RASP: A Drone-based Reconfigurable Actuation and Sensing Platform Towards Ambient Intelligent Systems

### v1

@@ Abstract.
Realizing consumer-grade drones that are as useful as robot vacuums throughout our homes or personal smartphones in our daily lives requires drones to sense, actuate, and respond to general scenarios that may arise. Towards this vision, we propose RASP, a modular and reconfigurable sensing and actuation platform that allows drones to autonomously swap onboard sensors and actuators in only $25$ seconds, allowing a single drone to quickly adapt to a diverse range of tasks. RASP consists of a mechanical layer to physically swap sensor modules, an electrical layer to maintain power and communication lines to the sensor/actuator, and a software layer to maintain a common interface between the drone and any sensor module in our platform. Leveraging recent advances in large language and visual language models, we further introduce the architecture, implementation, and real-world deployments of a personal assistant system utilizing RASP. We demonstrate that RASP can enable a diverse range of useful tasks in home, office, lab, and other indoor settings.
@@ 1. Introduction
Drones are experiencing increased utilization across numerous industries, including search and rescue, construction, delivery, agriculture, and many more (Wingtra, 2023; JOUAV, 2022; heliguy, 2020; Croptracker, 2019; Tucker, 2015; Johnson et al., 2021; Xia et al., 2023). The global commercial drone market is expected to grow to over 580 billion USD by 2030 (dro, 2023). While commercial drones are making significant impacts across various sectors, consumer drones (such as those from DJI (Shenzhen DJI Sciences and Technologies Ltd., 2022) or Skydio (Skydio, Inc., 2016)), that are more common in everyday life, tend to be confined to hobbyist or entertainment purposes, such as aerial photography and videography. However, we believe that the potential of drones in daily life as a versatile platform for mobile sensing and actuation extends far beyond these limited applications.
Imagine a drone in every home, much like robot vacuum cleaners, or each person having a drone acting as a personal “flying smartphone with interchangeable attachements” that you can direct through a variety of interfaces, such as voice commands. The utility of such a platform depends on the quantity and quality of tasks the drone is equipped to handle, which is heavily correlated with the number of sensing and actuation modalities it can support; the number of tasks, generality of a drone, and usefulness of a drone increases with each new sensing and actuation modality it supports.
We propose RASP, an adaptive, modular, and reconfigurable actuation and sensing platform that allows the same drone to adapt to diverse scenarios by quickly and efficiently switching out sensors and/or actuators depending on the situation. Figure 1 shows an overview of the workflow of RASP. The drone lands, communicates with the base station, and specifies the module to swap in. The base station secures the drone and actuates the conveyor belt and gripper to remove and replace the onboard sensor/actuator with the requested module. The unified module interface (UMI) ensures that the sensor/actuator module is firmly attached, connected, and can be accessed by the drone. Finally, the drone flies off with the new module for its new task.
@@
Our approach of modularizing a drone’s sensors and actuators offers a number of advantages compared to a static drone with all sensors and actuators attached. Much like single-lens reflex (DSLR) cameras with interchangeable lenses, 1) modularization allows for the creation of an ecosystem of sensors and actuators. This allows consumers to purchase only the sensors and actuators they need and improve drone reusability, expendability, and sustainability. A static drone may not have all the sensors and actuators you may need or want, which requires you to purchase a completely new drone, rather than just the sensor or actuator. 2) Layering and modularization enables you to evolve the drone, sensors, and actuators independently. New sensors and actuators purchased will still be compatible with your current drone as long as the interface remains the same and vice versa. 3) A drone carrying a single sensor or actuator can be designed much smaller, more agile, less noisy, and more suitable for closed environments, such as indoors.
We believe RASP to be one crucial step towards general drone platforms capable of aiding in many daily activities, much like the role of smartphones. To demonstrate this, we prototype and show how RASP could be integrated into a personal assistant system that also leverages static sensors in the environment (cameras) in conjunction with foundational models and penetrative AI (Xu et al., 2024) to satisfy a wide range of useful tasks in a home, lab, or office setting. For example, a drone could carry a temperature sensor to determine the warmest place to sit, before swapping in a payload module to carry a snack to an occupant. To the best of our knowledge, this deployment is the first end-to-end architectural, mechanical, and software design to autonomously control and reconfigure drones for versatile tasks leveraging foundational models. Our contributions are as follows.
@@
1. We create and implement RASP, a modular and reconfigurable sensing and actuation platform that allows a single drone to adapt to a wide range of scenarios. RASP allows drones to dynamically change their onboard sensing and actuation modules in under $25$ seconds.
@@
2. To realize RASP, we propose a three-layer architecture that physically replaces onboard modules, maintains data connection lines, and enables drones to communicate with sensing and actuation modules through a unified interface.
3. To show the importance of RASP for future ambient intelligent systems, we prototype and integrate RASP into a personal assistant system that takes advantage of static sensors in the environment (cameras), large language and vision language models, and the mobility of a drone to realize a wide range of useful applications in smart home, lab, office, and indoor settings.
@@ 2. Related Works
There are many reconfigurable and modular sensing platforms in existence. In addition to the platforms provided by open-source do-it-yourself (DIY) electronic vendors, such as Adafruit (Industries, 2022) and Sparkfun (Electronics, 2022), there are also platforms that operate on even lower resource microcontrollers, without an operating system, that are less flexible in the number of interfaces and configurations they support (Yi et al., 2018).  (Zhao et al., 2022, 2023) are reconfigurable sensing platforms based on the Raspberry Pi that have unified and generic hardware interfaces, allowing sensors to use the same set of connectors even if they are interfaced differently in software (e.g., UART, SPI, or I2C). Several platforms leverage the Berkeley TinyOS operating system (Levis et al., 2005), which allow developers to develop extremely long-lasting applications with great flexibility (Culler et al., 2002; O’Flynn et al., 2005; Edmonds et al., 2005; Harte et al., 2007; Mikhaylov et al., 2013; Mikhaylov and Huttunen, 2014). Despite the prevalence of reconfigurable and modular sensing platforms, there currently does not exist any platform that allow drones to reconfigure and adapt its onboard sensors and actuators automatically.
The idea of a smart home and other smart environments has existed for many decades. There are many smart home systems developed by research groups (Das et al., 2002; Mozer, 1998; Helal et al., 2005) and commercial products from companies like Google (LLC, 2016), Amazon (Amazon.com, 2016), and Apple (Inc., 2020). Most of these leverage static sensors and actuators at specific parts of the home to sense and respond to home or human activity. Though there are a few up and coming products (e.g., Ring Always Home Drone (rin, 2023)) and research works (Xia et al., 2021) that leverage drones for indoor and home scenarios, more research is needed in all areas (e.g., control, localization, perception, understanding) to realize a full drone-based system that can operate autonomously and take advantage of the sensing and compute from static sensors throughout indoor environments. (Xia et al., 2021) recently proposed a reconfigurable drone platform, but is only a demo abstract and leaves out many details on how it autonomously swaps modules.
Language and foundational models have seen a surge in usage and research due to their powerful capabilities in allowing computers to understand and interact using natural human modes of communication in an unprecedented manner. There are a few works that leverage language models as an interface for people to interact with their smart homes (King et al., 2023; Nie et al., 2022b, a) or drones and other robotics (Vemprala et al., 2023). However, to the best of our knowledge there are no works that attempt to realize a smart drone-based system for indoor settings that jointly utilizes the strengths of language models, static infrastructure, and a reconfigurable drone platform to perform a diverse range of useful tasks for occupants.
3. Modular and Reconfigurable Sensing and Actuation Drone Platform
@@
In this section, we introduce the key components of RASP, our reconfigurable and modular sensing and actuation platform and the drone system that it is built upon. Figure 2 shows the architecture of RASP. At a high level there are three major conceptual architectural components: the ground station or mechanical layer, electrical connection layer, and the software layer. These conceptual architectural components are implemented across five physical platform components: the ground station (Figures 3e and 5), sensor/actuation modules (Figures 3c-d and 4), unified module interface (UMI) (Figure 6), the carrier board (Figure 3b), and the physical drone (Figure 3a).
@@ 3.1. Architecture
Mechanical Layer. To enable reconfigurability, there needs to be a mechanism to physically swap out different sensor modules; the mechanical layer is responsible for this. The drone will land on the ground station, which will physically detach the current module onboard the drone and attach the newly requested module.
Electrical Connection Layer. Each sensor and actuation module needs to maintain data, communication, and power to the drone, despite being detachable; The electrical layer maintains these connections. We leverage magnets to accomplish this, maintaining physical connection between the module and the drone during a mission, while still allowing the mechanical layer to easily swap out modules.
Software Layer. In addition to maintaining physical and electrical connections, the platform needs to adapt to new modules on-the-fly even if they are swapped out. The software layer provides standardized protocols and libraries that allow the drone to discover, configure, control, and receive data from attached modules.
@@ 3.2. Platform Components
Drone platform. We designed the drone based on the open-source Crazyflie drone (cra, 2023), as shown in Figure 3. We chose the lightweight (27g) Crazyflie as our foundation due to its compact size, making it easier to operate in all environments, not just in open outdoor spaces. To increase its payload, we replaced its brushed motors (19000 Kv powered by a 1-cell battery) with more powerful brushless motors (3800 Kv powered by a 4-cell 850mAh battery). We also printed a protective guard, not only to protect its propellers, but to protect people and other objects from its propeller blades. The drone can hover for nearly 4 minutes, which is comparable to the Crazyflie drone.
@@
Figure 4. a) 3D structure attached to modules to allow the RASP to easily pick them up, b) and c) are displacement plots from stress analyses; the corresponding lowest factors of safety (FOS) are 1.180 and 4.664.
Sensor, Actuation, and Recharging Modules. RASP comes with a collection of sensor and actuation modules (full list in Table 1). These modules are attached or detached from the drone by the ground station/mechanical layer. The actuation module, shown in Figure 3d, is a container structure with a motor that opens and closes a hatch. Small items (e.g., medication, candy, pet food, etc.) can be loaded into this module for the drone to deliver.
@@
To make it easier for the ground station to pick up a module, we 3D printed and attached a structure shown in Figure 4. We performed force analysis (using SOLIDWORKS SimulationXpress) to select the optimal structure and thickness for the frame that is as light as possible to minimize extra weight while being sturdy enough to avoid being deformed when the gripper holds it. We measured that the grippers applies a 5 N force onto the sides of the frame when it grips the frame and a 10 N at the bottom of the frame when it brings the module down to the conveyor belt. To avoid bending at these forces, our force analysis showed when the frame is 1.5mm thick, it had a factor of safety (FOS) greater than 1 in both cases, indicating that the frame was able to withstand the forces applied. The mass of this frame is 11.97 g, which is negligible when compared to the rest of the drone (344.7 g).
@@
Ground station and automated takeoff and landing. The ground station primarily spans the mechanical layer, consisting of a gripper mechanism and a conveyor belt to position modules to swap in/out. We created the platform leveraging the chassis of the open-source Ender-3 3D printer (end, 2023) because it provides enough space for the drone to land, as well as enough mechanical control to remove onboard sensors and attach new sensors onto the drone.
We designed and benchmarked three different grippers with different slots and grooves and measured the offset from the center of the parked drone on the ground station. The optimal design, shown in Figure 3e, helps the module slide into a centered position even if the module is slightly offset because of errors in actuating the conveyor to move the module into place.
@@
Removing an attached sensor and ensuring that the new sensor is securely fastened requires the drone to land precisely at the specific location where the interchange occurs. However, precise drone landing and alignment is extremely difficult because nonlinearities in drone stability become more pronounced as the drone moves closer to the ground (ground effect) (Matus-Vargas et al., 2021). Rather than creating a flat landing pad that is similar to the landing pads for helicopters, we take inspiration from the Ring Always Home Drone, which uses a funnel (rin, 2023). As long as the drone lands within the larger radius of opening of the funnel, the drone will automatically slide towards the bottom of the funnel with an opening that latches onto the onboard module.
We test and benchmark two designs for the landing station, as shown in Figure 5c. One design is a normal funnel (Figure 5c-top). The second design contains grooves to help the drone slot into place (Figure 5c-bottom); the base of the drone also has protruding grooves that match the grooves in the funnel. For each design, we had the drone land and swap modules 75 times and found that the swap rates are fairly similar. However, the grooved design corrects for greater drone misalignments from the platform (higher maximum tolerable orientation and offset errors in Figure 5d). As such, we adopt the grooved design into RASP.
@@
Standardized 24-pin magnetic connections: Each sensor and actuation module in RASP comes equipped with two magnetic connectors each with 12 spring-loaded pins, and the carrier board has another set that matches. The magnets allow each module to remain attached and connected to the drone during flight, while being pliable enough to be easily removed by the mechanical layer if a module needs to be switched. The spring-loaded pins enables electrical connections to the sensors and actuators, while allowing the ground station to charge the drone.
Sensor and actuator drivers: The drone needs to support a variety of different sensors and actuators, each of which may be interfaced with differently. To support this, RASP implements a library of standard drivers, as part of the software layer, that support different communication and sampling schemes, namely: I2C, SPI, UART, GPIO, and ADC. We also provide developers with a driver template that allows them to easily incorporate new sensors and actuators into the RASP ecosystem.
Plug-and-play controller: Once a new module attaches to the carrier board, the drone needs to be able to start collecting data or using the actuator automatically, without needing to manually run any command. To enable this feature, we include a non-volatile memory chip on each module (M24C02 (STMicroelectronics, 2017) 2Kb I2C bus electrically erasable programmable read-only memory (EEPROM)), which stores the identity of the module. The carrier board will pull data from here every time a new module is attached and automatically downloads the corresponding driver from a repository.
@@
A drone-based personal assistant is one of the core ambient intelligent systems that we see RASP enabling. For several decades now, the number of smart devices, sensors, and actuators deployed throughout our homes, offices, and environments has been growing exponentially (Forecast, 2021). These static sensing devices can provide a coarse-grained and localized picture of events and conditions within the environment; obtaining a fine-grained high resolution picture would require dense deployments of sensors, covering every square inch of the environment, which is impractical. We believe that adapting RASP, drones, and other mobile robotics into smart environments and ecosystems could enable fine-grained sensing and actuation at the tip of our fingers without the need for dense static deployments.
Figure 7 shows the architecture and workflow of our prototype drone-based personal assistant. Static cameras in the environment provide a high-level and coarse-grained view of the environment; in future work and systems, we envision RASP being compatible with other static sensors, beyond cameras. When a user gives a voice command, the system leverages a visual-language model to analyze the scene, interpret the command, and output a list of potential locations of interest, along with the relevant sensor or actuator. Then, the system actuates the RASP drone with the sensor/actuator to each of the locations to complete the task.
For example, one task or question a person may want to know is “where is the warmest place to sit”? A standard RGB camera would not be able to accurately make this determination without additional sensors (e.g., temperature). The voice command and images are analyzed by a visual-language model, which outputs potential locations where a person could sit (e.g., chairs and couches). Then, the system outfits the RASP drone with a temperature module and actuates the drone to each of these locations to measure temperature. The time series data from the temperature sensor will then be analyzed by another language/foundational model to make the final determination. There are four classes of tasks that our system can accommodate.

### v2

# RASP: A Drone-based Reconfigurable Actuation and Sensing Platform for Engaging Physical Environments with Foundation Models

@@ Abstract.
Foundation models and large language models have shown immense human-like understanding and capabilities for generating text and digital media. However, foundation models that can freely sense, interact, and actuate the physical world like in the digital domain is far from being realized. This is due to a number of challenges including: 1) being constrained to the types of static devices and sensors deployed, 2) events often being localized to one part of a large space, and 3) requiring dense and deployments of devices to achieve full coverage. As a critical step towards enabling foundation models to successfully and freely interact with the physical environment, we propose RASP, a modular and reconfigurable sensing and actuation platform that allows drones to autonomously swap onboard sensors and actuators in only $25$ seconds, allowing a single drone to quickly adapt to a diverse range of tasks. We demonstrate through real smart home deployments that RASP enables FMs and LLMs to complete diverse tasks up to $85\%$ more successfully by allowing them to target specific areas with specific sensors and actuators on-the-fly.
@@ 1. Introduction
Foundation models (FM) and artificial general intelligence (AGI) have the potential to bring about massive technological and social change to the physical world by discovering, performing, and continually adapting to new applications and environments without needing to manually embed or program them. While LLMs and FMs have been shown to have human-like understanding and digital media generation capabilities, their ability to respond to events and actuate systems in the physical world are comparatively less explored. There are a number of works that incorporate language models to control physical robotic systems through voice (Vemprala et al., 2023; Nie et al., 2022b), from both industry and academia, such as the Figure01 AGI robot (Figure, 2024). The sensing and control in such scenarios are often localized to the vicinity of the autonomous agent.
Scaling up, there are few works that explore the use of LLMs to actuate our environments, particularly smart homes (King et al., 2023, 2024; Rivkin et al., 2023), where events and actions may occur anywhere in the space. These works generally focus on adapting LLMs as a human-like interface to actuate common internet-connected smart appliances (e.g., speakers, television, air conditioning, etc.). Much like how FMs enable general human language and sensory understanding and responses, our work explores how FMs could be used to enable a diverse range of general interactions with objects and spaces that likely cannot be actuated digitally through code. These interactions form a large portion of our daily interactions.
@@
For example, a person may want his/her home to physically bring a snack or medicine; this would require a robotic system that could physically identify and carry the payload. Another person, coming into an office area for the time may want to know the warmest desk to sit at in the building; this would require the building to utilize a dense deployment of temperature sensors. A third person in a chemical lab may want the building to monitor and notify him/her about the results of a chemical reaction that s/he started before leaving; this situation could likely be accomplished with a camera with special capabilities to detect these reactions. While it is possible to enable all these applications, these examples highlight two main challenges that prevent FMs and LLMs for achieving the same amount of autonomy in our physical environments as we have seen in the digital domain:
1. Each new application requires a new device or sensor, often with built-in special capabilities. Each of the three examples we discussed would require the user to purchase or engineer a system to satisfy that task. Evolving new applications and functionality in this way is also not scalable. Another way to state this challenge is that the generality and types of applications available are limited by the functionality of the devices and sensors present.
2. Events, tasks, and actions are often localized. Certain events and tasks a user may require the FM to analyze the entire space, but perform an action at a target location. For example, placing rat poison in areas around the home that will be most effective. If a specific sensing modality needed to make this judgment (e.g., moisture) is not present at a specific location, then the FM cannot obtain the information it needs to carry out the task. This leads to the third challenge.
3. Achieving full coverage across an entire space requires space-dependent dense deployments. The example of finding the “best desk” to sit requires a dense deployment of temperature sensors. While it is common for smart spaces to deploy smart devices and sensors, new applications may require dense deployments that need to be tailored to the layout of the environment, making it difficult to achieve full coverage for each new application.
Towards enabling general task execution by FMs in physical environments, we propose RASP, an adaptive, modular, and reconfigurable actuation and sensing drone platform that allows FMs and LLMs to adapt the same drone to diverse scenarios by efficiently switching out sensors and/or actuators depending on the situation. Figure 1 shows an overview of RASP. The drone lands, communicates with the base station, and specifies the module to swap in. The base station secures the drone and actuates the conveyor belt and gripper to remove and replace the onboard sensor/actuator with the requested module. The unified module interface (UMI) ensures that the sensor/actuator module is firmly attached, connected, and can be accessed by the drone. Finally, the drone flies off with the new module for its new task. RASP is an important step towards FMs and LLMs that can dynamically interact with occupants and the environment and addresses all of the challenges mentioned previously:
@@
1. Modularization enables easy integration of new applications. Much like single-lens reflex (DSLR) cameras with interchangeable lenses, modularization 1) allows for the creation of an ecosystem of sensors, actuators, and applications. Consumers can therefore purchase only the sensors and actuators for the applications they need and improve drone reusability, expendability, and sustainability. A static drone may not have all the required sensors and actuators, which would require purchasing a completely new drone. Layering and modularization also enables 2) evolving the drone and enabling new applications independently. New sensors and actuators purchased for new applications will still be compatible as long as the interface remains the same. Drones carrying a single sensor or actuator 3) can be designed much smaller, more agile, less noisy, and more suitable for closed environments, such as indoors.
2. Actuating a drone enables localized sensing and actuation. A reconfigurable drone platform enables FMs to dynamically specify what sensing or actuation modality at which location, without relying entirely on static sensors that may not have been deployed.
3. Drones can achieve full spatial coverage while allowing for sparser deployments. Because FMs can actuate a drone to any location, there no longer is a need for dense deployments for all types of sensors. Instead, a single drone can be used to service an entire space.
To demonstrate the utility of RASP for LLMs and FMs interfacing with physical environments, we prototype and show how RASP could be integrated into a personal assistant system that leverages static sensors deployed throughout the environment (cameras) in conjunction with foundation models and penetrative AI (Xu et al., 2024) to satisfy a wide range of useful tasks in a home, lab, or office setting. While there are several existing configurable drone platforms (Schiano et al., 2022; Derrouaoui et al., 2021; Perikleous et al., 2024; Hert et al., 2023), these works focus on the physical design and control of the drone rather than the sensing, analysis, and actuation capabilities. Moreover, these existing works require manual reconfiguration by humans, while RASP autonomously reconfigures the sensing and actuation capabilities on-the-fly. To the best of our knowledge, our work is the first to propose a drone platform with reconfigurable sensing to address challenges in coverage, localized sensing, and dense deployments for enabling general LLM and FM interactions with our physical environments. Our contributions are as follows.
@@
1. We propose and implement RASP, a modular and reconfigurable sensing and actuation platform that enables FMs to adapt the same drone to a wide range of scenarios through physical spaces. RASP equips FMs and LLMs with drones that can dynamically change their onboard sensing and actuation modules in under $25$ seconds.
@@
2. To realize RASP, we propose a three-layer architecture that physically replaces onboard modules, maintains data connection lines, and enables drones to communicate with sensing and actuation modules through a unified interface. Unlike existing reconfigurable drone platforms that focus on drone control and design, and require manual intervention to swap components, RASP focuses on systems and methods to autonomously reconfigure the drone’s onboard sensing and actuation capabilities, without human intervention.
3. We demonstrate how RASP can augment LLMs and FMs and easily enable new applications throughout our environments, beyond the capabilities of common IoT smart devices, by prototyping a personal assistant system that leverages both static sensors (cameras) and the mobility and flexibility of a reconfigurable drone. Our deployments in realistic scenarios demonstrate that RASP can improve the success rate of a diverse array of tasks by up to $85\%$.
@@ 2. Related Works
1. Language and foundation models. LLMs and FMs have seen a surge in usage and research due to their powerful capabilities in allowing computers to understand and interact using natural human modes of communication in an unprecedented manner. While most of these works focus on generating language, text, and digital media, there is a growing trend on adapting them for autonomous systems that interact with the physical world. Many such works target robotic platforms (Vemprala et al., 2023), including drones and robot vacuums, that only have an egocentric view of the vicinity around them. Works that leverage FMs and LLMs for interacting with larger spaces, namely smart homes, generally focus on actuating common smart appliances to better respond to the needs of occupants (King et al., 2023; Nie et al., 2022b, a). These works, leverage LLMs to create a natural language interface between humans and their environments, often leveraging internet-connected devices and targetting applications that are already widespread (e.g., television or air conditioning control). Our work focuses on enabling new applications and interactions that LLMs and FMs can have with occupants and their environments scalably without being limited to the constraints imposed by the sensors and functionalities implemented and hard coded into the devices found throughout the environment.
@@
2. Reconfigurable sensing and drone platforms. There are many reconfigurable and modular sensing platforms in existence. In addition to the platforms provided by open-source do-it-yourself (DIY) electronic vendors, such as Adafruit (Industries, 2022) and Sparkfun (Electronics, 2022), there are also platforms that operate on even lower resource microcontrollers, without an operating system, that are less flexible in the number of interfaces and configurations they support (Yi et al., 2018).  (Zhao et al., 2022, 2023) are reconfigurable sensing platforms based on the Raspberry Pi that have unified and generic hardware interfaces, allowing sensors to use the same set of connectors even if they are interfaced differently in software (e.g., UART, SPI, or I2C). Several platforms leverage the Berkeley TinyOS operating system (Levis et al., 2005), which allow developers to develop extremely long-lasting applications with great flexibility (Culler et al., 2002; O’Flynn et al., 2005; Edmonds et al., 2005; Harte et al., 2007; Mikhaylov et al., 2013; Mikhaylov and Huttunen, 2014).
There are a few reconfigurable drone platforms, but most focus on enabling flexible physical design and control of drones (Schiano et al., 2022; Derrouaoui et al., 2021; Perikleous et al., 2024; Hert et al., 2023). For example, (da Silva Ferreira et al., 2020) allows operators to reconfigure the number of rotors the drone with locking mechanisms that connect rotor modules. These works typically do not focus on sensing and actuation reconfiguration, unlike RASP. Additionally, they typically require a person to manually reconfigure the drone, while we introduce mechanisms that enables human-free configuration of the sensing and actuation. (Xia et al., 2021) recently proposed a reconfigurable drone platform, but is only a demo abstract and leaves out many details on how it autonomously swaps modules.
3. Modular and Reconfigurable Sensing and Actuation Drone
@@
Figure 2 shows the architecture of RASP. At a high level there are three major conceptual architectural components: the ground station or mechanical layer, electrical connection layer, and the software layer. These conceptual architectural components are implemented across five physical platform components: the ground station (Figures 3e and 4), sensor/actuation modules (Figures 3c-d), unified module interface (UMI) (Figure 5), the carrier board (Figure 3b), and the physical drone (Figure 3a).
@@ 3.1. Architecture
Mechanical Layer. On a functional level, the mechanical layer separates RASP from existing reconfigurable drone platforms by enabling the autonomous swapping of sensor and actuator modules. The drone lands on the ground station, which will physically detach the current module and attach the newly requested one. We propose novel swapping, including an augmented funnel shaped landing structure, to enable autonomous human-free swapping of modules as discussed in Section 3.2.
Electrical Connection Layer. This layer maintains the data, communication, and power connection between each sensor or actuation module with the drone, while being both detachable yet reliable mid mission. Existing reconfigurable drone platforms (da Silva Ferreira et al., 2020) generally incorporate slotted mechanical mechanisms. While these pins can maintain connection under great amounts of force, they are difficult to attach and remove, requiring a person to manually connect or disconnect. Attaching often requires precisely aligning slots or pins and applying acute pressure at junction points.
Instead, we leverage magnets to maintain connection. Attaching a magnet-based connector only requires moving the module close to the drone’s connector before the magnetic pull aligns all connections in place. Removing the module can be achieved without needing to apply acute force directly at the junction. Both of these characteristics allow magnet-based connectors to be more conducive for autonomous reconfigurable systems than purely mechanical connectors. We discuss in more detail in Section 3.2.
Software Layer. We propose standardized protocols and software libraries on top of the mechanical and electrical connection layers that allow the drone to discover, control, and receive data from attached modules on-the-fly.
@@ 3.2. Platform Components
Drone platform. We build RASP off the open-source Crazyflie drone (cra, 2023), as shown in Figure 3. We chose the lightweight (27g) Crazyflie as our foundation due to its compact size, making it easier to operate in all environments, especially indoors. To increase its payload, we replaced its brushed motors (19000 Kv powered by a 1-cell battery) with more powerful brushless motors (3800 Kv powered by a 4-cell 850mAh battery). We also printed a protective guard to protect its blades from objects and people. The drone can hover for nearly 4 minutes, which is comparable to the Crazyflie.
@@
Sensor, Actuation, and Recharging Modules. RASP comes with a collection of sensor and actuation modules (full list in Table 1). These modules are attached or detached from the drone by the ground station/mechanical layer and comes with a 3D printed structure that increases the surface for the ground station to pick up modules. The actuation module, shown in Figure 3d, is a container structure with a motor that opens and closes a hatch. Small items (e.g., medication, candy, pet food, etc.) can be loaded into this module for the drone to deliver.
@@
Ground station and automated takeoff and landing. The ground station primarily spans the mechanical layer, consisting of a gripper mechanism and a conveyor belt to position modules to swap in/out. We created the platform leveraging the chassis of the open-source Ender-3 3D printer (end, 2023) because it provides enough space for the drone to land, as well as enough mechanical control to remove onboard sensors and attach new sensors onto the drone. The gripper in Figure 3e removes and attaches modules onto the drone.
@@
Landing Platform. Removing an attached sensor and ensuring that the new sensor is securely fastened requires the drone to land precisely at the specific location where the interchange occurs. However, precise drone landing and alignment is extremely difficult because nonlinearities in drone stability become more pronounced as the drone moves closer to the ground (ground effect) (Matus-Vargas et al., 2021). Rather than creating a flat landing pad that is similar to the landing pads for helicopters, we take inspiration from the Ring Always Home Drone, which uses a funnel (rin, 2023). As long as the drone lands within the larger radius of opening of the funnel, the drone will automatically slide towards the bottom of the funnel with an opening that latches onto the onboard module (Figure 4c-top). To further improve alignment, we propose a new design that includes additional grooves to further improve module and drone alignment (Figure 4c-bottom).
We benchmark both the vanilla funnel and the grooved funnel landing platform design, as shown in Figure 4d. For each design, we had the drone land and swap modules 75 times and found that the swap rates are fairly similar. However, the grooved design corrects for greater drone misalignments from the platform, as reflected in the higher maximum tolerable orientation and offset errors. As such, we adopt our proposed grooved funnel design into RASP.
@@
Standardized 24-pin magnetic connections: Unlike existing reconfigurable drone and sensing platforms that use mechanical connectors, each sensor and actuation module in RASP comes equipped with two magnetic connectors each with 12 spring-loaded pins, with corresponding connectors on the carrier board. The magnets allow each module to remain attached and connected to the drone during flight, while being pliable enough to be easily removed by the mechanical layer. While mechanical pins and slots are the standard for maintaining stable connections, they are difficult to attach and remove without human intervention. Magnets allow attaching and removing modules with much less precision, with enough force to maintain connection mid-mission. Magnetic connections only require placing the module in the vicinity of the drone to attach the module and diffused force to remove, compared to standard mechanical connectors.
Sensor and actuator drivers: To support a variety of sensors and actuators that may each be interfaced differently, RASP implements a library of standard drivers, as part of the software layer, that support different communication and sampling schemes, namely: I2C, SPI, UART, GPIO, and ADC. We also provide developers with a driver template that allows them to easily incorporate new sensors and actuators into the RASP ecosystem.
Plug-and-play controller: Once a new module attaches to the carrier board, the drone needs to be able to start collecting data or using the actuator automatically, without needing to manually run any command. To enable this feature, we include a non-volatile memory chip on each module (M24C02 (STMicroelectronics, 2017) 2Kb I2C bus electrically erasable programmable read-only memory (EEPROM)), which stores the identity of the module. The carrier board will pull data from here every time a new module is attached and automatically downloads the corresponding driver.
@@
In this section, we detail the prototype of RASP into a personal assistant application that has information about the space through a deployment of static sensors (namely cameras). Sections 4.1 and 4.2 introduce methods for adapting visual language models to analyze and understand sensor data from the environment, while Sections 4.4 and 4.3 introduce how FMs can integrate RASP into their pipeline.
Figure 6 shows the workflow of our prototype. Static cameras in the environment provide a high-level and coarse-grained view of the environment; in future work and systems, we envision RASP being compatible with other static sensors, beyond cameras. When a user gives a voice command, the system leverages a visual-language model to analyze the scene, interpret the command, and output a list of potential locations of interest, along with the relevant sensor or actuator. Then, the system actuates the RASP drone with the sensor/actuator to each of the locations to complete the task.
For example, one task or question a person may want to know is “where is the warmest place to sit”? A standard RGB camera would not be able to accurately make this determination without additional sensors (e.g., temperature). The voice command and images are analyzed by a visual-language model, which outputs potential locations where a person could sit (e.g., chairs and couches). Then, the system outfits the RASP drone with a temperature module and actuates the drone to each of these locations to measure temperature. The time series data from the temperature sensor will then be analyzed by another language/foundational model to make the final determination. There are four classes of tasks that we explore in this deployment, multiple of which are currently not commonly packaged into existing IoT smart devices: